# Magnetic phase diagram of multiferroic and magnetocaloric $\mathbf{TmFeO_{3}}$

### v2

@@ Abstract
Neutron diffraction experiments of $\rm{TmFeO_{3}}$ single crystals were performed in the external magnetic fields. The field along c-axis increases temperature of spin-reorientation transition $T_{SR}$ from phase $\Gamma 4$ to $\Gamma 2$. Application of the field along b-axis led to the decrease of $T_{SR}$ and to the formation of new phases. Based on the temperature and field dependence of the Bragg reflection intensity, the configuration of magnetically induced phases was proposed. The magneto-structural effects were observed.
@@ I Introduction
As it is known from a number of previous works on orthoferrites (e.g. [17, 18]), the exchange interaction within the Fe-sublattice plays the main role in the formation of the magnetic structure below TN. For $\rm{TmFeO_{3}}$ the strongest interaction is the superexchange one through O between the nearest neighbors in the Fe-chains along the b-axis and within the ac-plane $J_{b}^{\rm{Fe}}=5.15$ meV and $J_{ac}^{\rm{Fe}}=4.74$ meV respectively [17]. The exchange field from ordered $\rm{Fe^{3+}}$ subsystem induces an ordered magnetic moment on $\rm{Tm^{3+}}$ ions, which is very weak at the temperature above $T_{SR}$, but becomes more noticeable below spin-reorientation transition [19]. Thus, when the temperature is decreased, the exchange between Fe and Tm sublattices, $J^{\rm{Fe-Tm}}$ becomes effective, which leads to spin reorientation transition.
@@ II Experimental part
High quality single crystals of $\rm{TmFeO_{3}}$ were grown by the flux method [20] using the optical floating zone technique (FZ-4000, Crystal Systems Corporation) with the natural isotope mixture. X-Ray studies were performed at the Rigaku SmartLab diffractometer at the Petersburg Nuclear Physics Institute (PNPI). These measurements were performed with wavelength 1.54 Å, $2\theta$ range $20^{\circ}-135^{\circ}$ and the temperature range $T=50-300$ K using powder obtained by grinding of the crystal. Refinement of crystal structure was made using the FullProf Suite software package [21]. Single crystal with cylindrical shape and dimensions of $4\times 4\times 8$ mm3 was used for the neutron studies. The neutron diffraction experiments were performed at the Institut Laue-Langevin. The orientation of the crystal was obtained by OrientExpress [22]. This is an automated Laue neutron diffractometer based upon two high-performance image-enhanced CCD cameras coupled to a large-area neutron scintillator and allows fast orientation of the crystal. Magnetic structure studies were performed on diffractometer D23 [23] in external magnetic fields with wavelength 1.28 Å. Two sets of measurements were made for the direction of magnetic fields along crystal axes b and c with constant field values. For both sample orientations, sets of selected peaks were measured for each combination of temperature 2, 20, 40, 60, 75, 88, 100 K and magnetic field 0, 0.5, 1, 2, 3, 4, 5, 5.8 T. In addition, for both sample orientations, we measured intensities of selected peaks at temperatures 80, 90, 95, 105, and 110 K without magnetic field; this temperature interval includes the temperature range of Fe spin reorientation. Furthermore, two sets of all available reflections (about 500) at temperature $T=20$ K were collected for two different sample orientations. The program Mag2Pol [24] was used then to calculate and refine the magnetic structure.
@@ III.1 Structure study
Our X-ray studies confirmed that the crystal structure of $\rm{TmFeO_{3}}$ is described by the space group Pnma #62 IT [25] at all temperatures of measurement. In this structure Fe-ions occupy position 4b, Tm – 4c, O1 and O2 – 4c and 8d respectively. Obtained parameters of the crystal structure of $\rm{TmFeO_{3}}$ at different temperatures are shown in Table 1 and they agree well with those ones in works [26] [27]. The unit cell parameter c increases monotonically with increasing temperature. The unit cell size along a and b decreases with increasing temperature up to the spin reorientation transition $T_{SR}$ and increases monotonically above $T_{SR}$, which was also observed in [28, 29].
@@
[…]
The magnetic structure was refined simultaneously for two datasets collected with different crystal orientations at $T=20$ K without magnetic field. In magnetic neutron scattering, the magnetic form factor $F(q)$ defines the $q$-dependence of the magnetic scattering amplitude of a single ion in a such way that the magnetic signal at large $q$ – large angles – is negligible in comparison with the nuclear one. Since we have no dataset above Néel temperature, in the first stage of fitting, reflections measured at theta angles greater than 60 degrees were selected from the collected dataset and the scale factor and extinction parameters were fitted. In the second stage of fitting, the remaining reflections, collected at low angles, were used to estimate the values of Fe and Tm magnetic moments. Then magnetic moments values were obtained to be equal $M_{x}^{Fe}=0$, $M_{y}^{Fe}=4.93(22)$, $M_{z}^{Fe}=-0.51(9)$ and $M_{x}^{Tm}=0$, $M_{y}^{Tm}=0$, $M_{z}^{Tm}=0.37(8)$ $\mu_{B}$ with R-factor = 11.2. The refinement with free parameters for $M_{x}^{Fe}$ gives zero for these values within esds and $M_{y}^{Tm}$ zero due to the symmetry restrictions, then these parameters were fixed equal to zero. Magnetic ordering of the $\rm{Tm^{3+}}$ sublattice corresponds to the irreducible representation $\Gamma 2$. The obtained magnetic moments values are in good agreement with those ones, obtained for $\rm{TmFeO_{3}}$ in work [35]. Figure 1b shows the obtained magnetic structure at $T=20$ K without magnetic field.
@@
Table 2: Correspondence of (anti)ferroelectric and magnetic modes in Pnma crystal $\rm{TmFeO_{3}}$. Lower part of the Table show which ($hkl$) reflections gives intensity of each magnetic mode. ”e” and ”o” symbols denote even or odd parity of Miller indices, respectively.
@@ F A C G
$\mathbf{d_{2}}$ $\mathbf{S_{2}}$ $\left(0,\frac{1}{2},\frac{1}{2}\right)$ $\left(-x,\frac{3}{4},-z\right)$ + -- + --
$\mathbf{d_{3}}$ $\mathbf{S_{3}}$ $\left(\frac{1}{2},\frac{1}{2},0\right)$ $\left(-x+\frac{1}{2},\frac{3}{4},z+\frac{1}{2}\right)$ + -- -- +
$\mathbf{d_{4}}$ $\mathbf{S_{4}}$ $\left(\frac{1}{2},0,0\right)$ $\left(x+\frac{1}{2},\frac{1}{4},-z+\frac{1}{2}\right)$ + + -- --
@@
Table 3: Irreducible representations (irreps) of the space group Pnma built based on the $\mathbf{k}$-vector equal to ($0,0,0$) with corresponding magnetic space groups (MSG) and allowed magnetic modes for (4b) and (4c) Wyckoff positions.
@@
Figure 1: $\rm{TmFeO_{3}}$ structure. Red spheres are Tm ions, blue – Fe, yellow – O. Yellow lines show oxygen octahedrons around iron ions. a) Crystal structure of $\rm{TmFeO_{3}}$ at $T=300$ K; b) Magnetic structure of $\rm{TmFeO_{3}}$ at $T=20$ K. Blue arrows – direction of Fe magnetic moment, red arrows – Tm magnetic moments, the magnitude of the arrow for Tm is increased three times for convenience of demonstration; c) Antiferroelectric order of $\rm{TmFeO_{3}}$, green arrows – electric dipole moment.
@@ III.2 Magnetic scattering dependence on magnetic field
Figure 2 shows temperature dependencies two of them: reflections $(-110)$ (G-type) and $(-220)$ (F-type). The intensity fall at the zero field dependence of the G-type reflection in the temperature range of 83 – 93 K (Figure 2a) corresponds to the spin-reorientation transition from $\Gamma 4$ to $\Gamma 2$ phase through mixed phase $\Gamma 24$ with the rotation of the Fe main antiferromagnetic G-type order from the c-axis to the b-axis. The effective magnetic moment in that case decreases about twice, that leads to such intensity changes. The small increase in intensity below $T\approx 40$ K is probably associated with the additional insert of the induced magnetic moment on $\rm{Tm^{3+}}$ ions. The F-type reflection $(-220)$ keeps its intensity at zero field temperature dependence as expected.
The application of a magnetic field in two different directions has a different effect on the temperature dependence of Bragg reflections. At the field along b-axis, intensity of G-type reflection $(-110)$ goes to lower values in much smoother manner then it was without field, reaching lowest values at lower temperatures. It can also be seen that these lowest values are higher than that one which was observed without the field. And, at some temperatures – below 20 K for 1 T, below 40 K for 2 – 4 T – the intensity of $(-110)$ increases considerably. The similar situation takes place for the F-type reflection $(-220)$ that is increase of intensity at low temperatures when field was applied along b-axis (Figure 2b). The temperature behavior of these reflections at field 5.8 T looks very smooth and its intensity for $(-110)$ does not reach such a high value, as it was observed for smaller fields.
Figure 2: Temperature dependencies of reflections intensities ($-110$), ($-220$) for external magnetic field directed a) b) along the b-axis and c) d) along the c-axis.
@@
At the low temperature and high magnetic fields we have also observed significant increase of intensity of A-type peaks ($-410$), ($-210$), ($210$), moderate increase of intensity of ($-111$) and small/negligible change of intensity of ($113$), ($012$) reflections. These changes are probably connected with the contribution from rare earth subsystem. This behavior is indirectly correlated with the data of magnetic entropy change $\Delta S$, where the maximum of $\Delta S$ lies at 10 K in fields up to 3 T, and then shifts to the 15 – 20 K temperature range at fields over 3 T [16].
Below 40 K and above 3 T we also observed peak shifts in the omega angle from the centre of the scan window determined at 20 K without magnetic field. We interpret this as a magnetostriction effect. Interestingly, the amplitude of this shift strongly depends on the $hkl$ indices of reflections and varies for two magnetic field directions. For the magnetic field along the c-axis we observed positive shifts for all investigated reflections with significantly higher shifts for F$(1,0,1)$, A(1,3,1)1,3,1)1 , 3 , 1 ), and C$(0,2,1)$ peaks. For the magnetic field along the b-axis, peaks with positive $h$ and $k$ show positive shift, while peaks with negative $h$ and positive $k$ show negative shift. Similar strong magneto-structural coupling in $\rm{TmFeO_{3}}$ was reported in study [39]. It could be supposed to be a manifestation of symmetry lowering, and this should be checked in the following studies.

### v5

# Magnetic phase diagram of magnetocaloric $\mathbf{TmFeO_{3}}$

@@ Abstract
Neutron diffraction experiments of $\rm{TmFeO_{3}}$ single crystals were performed in the external magnetic fields. The field along c-axis increases temperature of spin-reorientation transition $T_{SR}$ from phase $\Gamma 4$ to $\Gamma 2$. Application of the field along b-axis led to the decrease of $T_{SR}$ and to the formation of new phases. Based on the temperature and field dependence of the Bragg reflection intensity, the configuration of magnetically induced phases was proposed.
@@ I Introduction
As it is known from a number of previous works on orthoferrites (e.g. [17, 18]), the exchange interaction within the Fe-sublattice plays the main role in the formation of the magnetic structure below $T_{N}$. For $\rm{TmFeO_{3}}$ the strongest interaction is the superexchange one through O between the nearest neighbors in the Fe-chains along the b-axis and within the ac-plane $J_{b}^{\rm{Fe}}=5.15$ meV and $J_{ac}^{\rm{Fe}}=4.74$ meV respectively [17]. The exchange field from ordered $\rm{Fe^{3+}}$ subsystem induces an ordered magnetic moment on $\rm{Tm^{3+}}$ ions, which is very weak at the temperature above $T_{SR}$, but becomes more noticeable below spin-reorientation transition [19]. Thus, when the temperature is decreased, the exchange between Fe and Tm sublattices, $J^{\rm{Fe-Tm}}$ becomes effective, which leads to spin reorientation transition.
@@ II Experimental part
High quality single crystals of $\rm{TmFeO_{3}}$ were grown by the flux method [20] using the optical floating zone technique (FZ-4000, Crystal Systems Corporation) with the natural isotope mixture. X-Ray studies were performed at the Rigaku SmartLab diffractometer at the Petersburg Nuclear Physics Institute (PNPI). These measurements were performed with wavelength 1.54 Å, $2\theta$ range $20^{\circ}-135^{\circ}$ and the temperature range T=50300T=50-300 K using powder obtained by grinding of the crystal. Refinement of crystal structure was made using the FullProf Suite software package [21]. Single crystal with cylindrical shape and dimensions of $4\times 4\times 8$ mm3 was used for the neutron studies. The neutron diffraction experiments were performed at the Institut Laue-Langevin. The orientation of the crystal was obtained by OrientExpress [22]. This is an automated Laue neutron diffractometer based upon two high-performance image-enhanced CCD cameras coupled to a large-area neutron scintillator and allows fast orientation of the crystal. Magnetic structure studies were performed on diffractometer D23 [23] in external magnetic fields with wavelength 1.28 Å. Two sets of measurements were made for the direction of magnetic fields along crystal axes b and c with constant field values. For both sample orientations, sets of selected peaks were measured for each combination at temperatures 2, 20, 40, 60, 75, 88, 100 K and magnetic fields 0, 0.5, 1, 2, 3, 4, 5, 5.8 T. In addition, for both sample orientations, we measured intensities of selected peaks at temperatures 80, 90, 95, 105, and 110 K without magnetic field; this temperature interval includes the temperature range of Fe spin reorientation. Furthermore, two sets of all available reflections (about 500) at temperature $T=20$ K were collected for two different sample orientations. The data that support the findings of this article are openly available [24]. The program Mag2Pol [25] was used then to calculate and refine the magnetic structure.
@@ III.1 Structure study
Our X-ray studies confirmed that the crystal structure of $\rm{TmFeO_{3}}$ is described by the space group Pnma #62 IT [26] at all temperatures of measurement. In this structure Fe-ions occupy position 4b, Tm – 4c, O1 and O2 – 4c and 8d respectively. Obtained parameters of the crystal structure of $\rm{TmFeO_{3}}$ at different temperatures are shown in Table 1 and they agree well with those ones in works [27] [28]. The unit cell parameters increase monotonically with increasing temperature, that was also observed in  [29, 30].
@@
where G describes the main antiferromagnetic component of the magnetic structure, F is the ferromagnetic vector, weak antiferromagnetic components C and A describe the weak canting of the magnetic moments. It is convenient situation to detect the mode of magnetic alignment of elements at 4b site in the crystal with space group Pnma. In this case different magnetic modes give contributions to Bragg reflections with different parity: mode A gives a contribution to reflections with $h+l$ even, $k$ odd; Ckk even, $h+l$ odd; Fh+lh+l even, $k$ even; Gh+lh+l odd, $k$ odd (see Table 2). $\mathbf{S_{1},S_{2},S_{3},S_{4}}$ means projections of Fe magnetic moments on arbitrary axis. In case $\rm{TmFeO_{3}}$, Fe-magnetic sublattice is ordered below temperature $T_{N}=635$ K in a G-type structure $\Gamma 4$ ($A_{x},F_{y},G_{z}$). Irreducible representation $\Gamma 4$ of the space group Pnma is build based on the propagation vector $\textbf{k}=(0,0,0)$ and it is equivalent to Pn’ma’ magnetic space group (see Table 3) with small spin canting caused by the Dzyaloshinskii-Moriya (DM) interaction that induces a weak ferromagnetic (FM) moment. Spin-reorientation transition begins at decreasing temperature below $T_{SR1}=93$ K. This transition goes through the mixed phase $\Gamma 24$ and at temperature $T_{SR2}=83$ K the main G-type component of Fe-spins lies along the b-axis and the magnetic phase $\Gamma 2\ (C_{x},G_{y},F_{z})$ is forming in the compound. Symmetry analysis [31] shows that continuous spin reorientation requires symmetry lowering at least to a monoclinic one. This statement is true only for continuous spin reorientation (i.e. for one magnetic domain where magnetic moment continuously changes direction with continuous change of temperature), while spin reorientation in $\rm{TmFeO_{3}}$ has been reported in the form of changing the population of magnetic domains [32] – such mechanism of spin reorientation is compatible with orthorhombic symmetry. In our studies, we have not observed lowering orthorhombic symmetry of lattice in the spin reorientation region that is also in the agreement with [33, 34]. As discussed in [19, 35] both modes below and above the spin reorientation transition correspond to the same ’AFB’ set of directions of magnetic moment. Detailed crystallographic and magnetic symmetry analysis of rare-earth orthoferrites was presented in [19] and analysis extension to (anti)ferroelectric and toroidal moments ordering was discussed in [35].
The magnetic structure was refined simultaneously for two datasets collected with different crystal orientations at $T=20$ K without magnetic field. In magnetic neutron scattering, the magnetic form factor $F(q)$ defines the $q$-dependence of the magnetic scattering amplitude of a single ion in a such way that the magnetic signal at large $q$ – large angles – is negligible in comparison with the nuclear one. Since we have no dataset above Néel temperature, in the first stage of fitting, reflections measured at $|Q|>8.5$ r.l.u. (two theta angles greater than 60 degrees) were selected for the structure refinement from the collected dataset and the scale factor and extinction parameters were fitted. In the second stage of fitting, the remaining reflections, collected at low angles, were used to estimate the values of Fe and Tm magnetic moments. Neutron paths in the sample were close each other in magnitude for reflections with different hklhkl, therefore absorption correction did not affect a lot on the results of the refinement. Then magnetic moments values were obtained to be equal $M_{x}^{Fe}=0$, $M_{y}^{Fe}=4.93(22)$, $M_{z}^{Fe}=-0.51(9)$ and $M_{x}^{Tm}=0$, $M_{y}^{Tm}=0$, $M_{z}^{Tm}=0.37(8)$ $\mu_{B}$ with R-factor = 11.2. The refinement with free parameters for $M_{x}^{Fe}$ gives zero for these values within esds and $M_{y}^{Tm}$ zero due to the symmetry restrictions, then these parameters were fixed equal to zero. Magnetic ordering of the $\rm{Tm^{3+}}$ sublattice corresponds to the irreducible representation $\Gamma 2$. The obtained magnetic moments values are in good agreement with those ones, obtained for $\rm{TmFeO_{3}}$ in work [36]. Figure 1b shows the obtained magnetic structure at $T=20$ K without magnetic field.
@@
Table 2: Correspondence of (anti)ferroelectric and magnetic modes in Pnma crystal $\rm{TmFeO_{3}}$. Lower part of the Table show which (hklhkl) reflections gives intensity of each magnetic mode. ”e” and ”o” symbols denote even or odd parity of Miller indices, respectively.
@@ F A C G
$\mathbf{d_{2}}$ $\mathbf{S_{2}}$ $\left(0,\frac{1}{2},\frac{1}{2}\right)$ $\left(-x,\frac{3}{4},-z\right)$ + - + -
$\mathbf{d_{3}}$ $\mathbf{S_{3}}$ $\left(\frac{1}{2},\frac{1}{2},0\right)$ $\left(-x+\frac{1}{2},\frac{3}{4},z+\frac{1}{2}\right)$ + - - +
$\mathbf{d_{4}}$ $\mathbf{S_{4}}$ $\left(\frac{1}{2},0,0\right)$ $\left(x+\frac{1}{2},\frac{1}{4},-z+\frac{1}{2}\right)$ + + - -
@@
Table 3: Irreducible representations (irreps) of the space group Pnma built based on the $\mathbf{k}$-vector equal to ($0,0,0$) with corresponding magnetic space groups (MSG) and allowed magnetic modes for 4b and 4c Wyckoff positions.
@@
Figure 1: (a) Crystal structure of $\rm{TmFeO_{3}}$ at $T=300$ K; (b) Magnetic structure of $\rm{TmFeO_{3}}$ at $T=20$ K. Blue arrows – Fe magnetic moments, red arrows – Tm magnetic moments, the magnitude of the arrow for Tm is increased five times for the sake of visualization; c) Antiferroelectric order of $\rm{TmFeO_{3}}$, green arrows – electric dipole moment.
@@ III.2 Magnetic scattering dependence on magnetic field
Figure 2 shows temperature dependencies two of them: reflections (110)(-110) (G-type) and (220)(-220) (F-type). The intensity fall at the zero field dependence of the G-type reflection in the temperature range of 83 – 93 K (Figure 2a) corresponds to the spin-reorientation transition from $\Gamma 4$ to $\Gamma 2$ phase through mixed phase $\Gamma 24$ with the rotation of the Fe main antiferromagnetic G-type order from the c-axis to the b-axis. The effective magnetic moment in that case decreases about twice, that leads to such intensity changes. The small increase in intensity below $T\approx 40$ K is probably associated with the additional insert of the induced magnetic moment on $\rm{Tm^{3+}}$ ions. The F-type reflection (220)(-220) keeps its intensity at zero field temperature dependence as expected.
The application of a magnetic field in two different directions has a different effect on the temperature dependence of Bragg reflections. At the field along b-axis, intensity of G-type reflection (110)(-110) goes to lower values in much smoother manner then it was without field, reaching lowest values at lower temperatures. It can also be seen that these lowest values are higher than that one which was observed without the field. And, at some temperatures – below 20 K for 1 T, below 40 K for 2 – 4 T – the intensity of (110)(-110) increases considerably. The similar situation takes place for the F-type reflection (220)(-220) that is increase of intensity at low temperatures when field was applied along b-axis (Figure 2b). The temperature behavior of these reflections at field 5.8 T looks very smooth and its intensity for (110)(-110) does not reach such a high value, as it was observed for smaller fields.
Figure 2: Temperature dependencies of reflections intensities (110-110), (220-220) for external magnetic field directed a) b) along the b-axis and c) d) along the c-axis.
@@
At the low temperature and high magnetic fields we have also observed significant increase of intensity of A-type peaks (410-410), (210-210), (210210), moderate increase of intensity of (111-111) and small/negligible change of intensity of ($113$), (012012) reflections. These changes are connected with the contribution from rare earth subsystem, as could be seen from following refinements. This behavior is indirectly correlated with the data of magnetic entropy change $\Delta S$, where the maximum of $\Delta S$ lies at 10 K in fields up to 3 T, and then shifts to the 15 – 20 K temperature range at fields over 3 T [16].
The sets of some tens reflections collected at temperatures 2, 20, 40, 60, 75, 88, 100 K and magnetic fields along 0, 0.5, 1, 2, 3, 4, 5, 5.8 T permitted to perform the refinement of magnetic structures realized at different condition (see Table 4). The accuracy of the magnetic moment determination is defined by the comparatively low number of measured peaks. One could mark, that in the $\Gamma 4$ phase magnetic moments on $\rm{Fe^{3+}}$ ions do not change their values under the applied field keeping $M_{z}^{Fe}=4.7(2)\ \mu_{B}$ for $G_{z}$ and $M_{y}^{Fe}=0.3(1)\ \mu_{B}$ for $F_{y}$. While $F_{y}$-component of $Tm^{3+}$ ions in $\Gamma 4$ demonstrate linear dependence on the applied field, thus changing from $M_{y}^{Tm}=0.5(2)\ \mu_{B}$ at zero field up to $M_{y}^{Tm}=1.3(3)\ \mu_{B}$ at 5.8 T field along b-axis. A different situation for that field direction takes place in $\Gamma 2$ phase. In this phase all magnetic components keep their values independent on external field. The refinement of the mixed phase $\Gamma 24$ provides confirmation, that this phase consists of the mixture of two magnetic domains with configuration $\Gamma 4$ and $\Gamma 2$ that is consistent with orthorhombic symmetry. For the temperature range 2 – 30 K and fields 2 – 4 T along b-axis, the refinement gives for $\rm{Fe^{3+}}$ subsystem antiferromagnetic alignment along c-axis of type $G_{z}$ and ferromagnetic alignment of type, that is characteristic of $\Gamma 4$ configuration. At the same time, for $\rm{Tm^{3+}}$ subsystem, the refinement gives non-zero $C_{x}$ and $F_{z}$ components, that corresponds to $\Gamma 2$ configuration. For magnetic fields 5 – 5.8 T at the temperatures 2 – 30 K we obtained $F_{y}$ and $G_{z}$ components for $\rm{Fe^{3+}}$, and non-zero $F_{y}$-component for $\rm{Tm^{3+}}$ subsystem, which depending from magnetic field has values $M_{y}^{Tm}=-0.5(3)--1.0(5)\ \mu_{B}$.
Table 4: Magnetic structures realized in $\rm{TmFeO_{3}}$ at different temperatures and magnetic fields. Roman numbers in column ”Configuration” correspond to spaces at phase diagram (Fig. 3). The specified values of magnetic moments are the average values for each configuration.
Configuration Irrep Mode $M_{Fe},\ \mu_{B}$ $M_{Tm},\ \mu_{B}$
I $\Gamma 4$ $A_{x}$
$F_{y}$ 0.3(1) 0.5(2) – 1.3(2)
$G_{z}$ 4.7(2)
II $\Gamma 2$ $C_{x}$
$G_{y}$ 4.9(2)
$F_{z}$ -0.4(1) 0.4(1)
$\Gamma 4$ $A_{x}$
$F_{y}$ 0.3(1) 0.5 – 1.3(2)
$G_{z}$ 4.8(2)
III $\Gamma 2$ $C_{x}$
$G_{y}$ 4.9(2)
$F_{z}$ -0.4(1) 0.4(1)
IV $\Gamma^{Fe}_{4}$ $A_{x}$
$F_{y}$ 0.3(1)
$G_{z}$ 4.8(3)
$\Gamma^{Tm}_{2}$ $C_{x}$ 0.3(1)
$G_{y}$
$F_{z}$ 0.4(1)
V $\Gamma 4$ $A_{x}$
$F_{y}$ 0.3(1) -0.5(3) – -1.0(3)
$G_{z}$ 4.7(3)